OpenStudy (anonymous):

You roll two dice, what is the probability of getting a sum of 5, given the first die is a three?

OpenStudy (kropot72):

The question reduces to finding the probability of rolling a 2 on a die.

OpenStudy (anonymous):

So that would be 2/36 where there are 36 possibilities?

OpenStudy (anonymous):

 This should help

OpenStudy (kropot72):

Not really. It is given that the first die is a three. The number rolled on the first die has no effect on the number rolled on the second die. To make a sum of 5, we need a 2 on the second die. The probability of rolling a 1, 2, 3, 4, 5 or 6 is 1/6.

OpenStudy (anonymous):

Ahh! Thank you both, the graph made sense and kropot72 cleared it up. So, it is 1/6.

OpenStudy (anonymous):

It's actually (1/6)^2

OpenStudy (anonymous):

The 2 represents the two dice?

OpenStudy (kropot72):

Yes, 1/6 is correct. (1/6)^2 ignores that it is given that the first die is a 3. (1/6)^2 is the probability of rolling a 3 on the first die and a 2 on the second die.

OpenStudy (anonymous):

Ah, my mistake.

OpenStudy (kropot72):

np :)

OpenStudy (anonymous):

I understood that 5/36 reduces down to 1/6. Thank you both very much!

OpenStudy (kropot72):

Note that (1/6)^2 stands for \[(\frac{1}{6})^{2}=\frac{1}{6}\times\frac{1}{6}=\frac{1}{36}\] and is not the correct answer.

OpenStudy (anonymous):

Yes, 1/6 and not 1/6^2.

OpenStudy (kropot72):

Correct :)